OpenStudy (anonymous):

The width and length of a rectangle (in feet) are consecutive odd integers. If the length is increased by 5 feet, the area of the resulting rectangle is 60 square feet. What is the area of the original rectangle? 25 ft2 30 ft2 35 ft2

OpenStudy (callisto):

Let x be the original width, then the original length = x+2 New length = (x+2)+5 = x+7 Area = length x width 60 = (x+7) x x^2 + 7x -60 =0 x=5 original area = x (x+2) = 5(5+2) =? Can you work it out?

OpenStudy (anonymous):

doesn't it come to 30ft2

OpenStudy (callisto):

Why does it?

OpenStudy (anonymous):

35ft2